OpenStudy (anonymous):

integral of 2 cos (theta/2)??? I dont really get why it is 4sin(theta/2)...

OpenStudy (anonymous):

@mathstudent55

OpenStudy (sidsiddhartha):

this is a standard integration \[\int\limits_{}^{}\cos(ax)dx=\frac{ 1 }{ a }\sin(ax)+c\]

OpenStudy (anonymous):

wouldnt that be just sin(theta/2)? because the 1/2 cancels...

OpenStudy (sidsiddhartha):

so.... \[\int\limits_{}^{}2 \cos(\frac{ \theta }{ 2 })d \theta=2\int\limits_{}^{}\cos \frac{ \theta }{ 2 }d \theta\\=2*\frac{ \sin \frac{ \theta }{ 2 } }{ 1/2 }+c\] ok?

OpenStudy (sidsiddhartha):

got this?

OpenStudy (anonymous):

ok, i think that makes sense

OpenStudy (anonymous):

yea, I think so

OpenStudy (sidsiddhartha):

good :)